OpenStudy (anonymous):

What number must you add to complete the square? x^2 + 4x = 7

OpenStudy (anonymous):

x^2+2(2)(x)+2^2 - 2^2= 7

OpenStudy (anonymous):

what do i add to complete the square ? what #?

OpenStudy (anonymous):

@annas when you put it like that, it's hard to understand for me as well lol.

OpenStudy (anonymous):

are you sure

OpenStudy (anonymous):

\[\pm 4\]

OpenStudy (anonymous):

\[x^2 + 4x + 4 = 7 + 4\] better version of anna

OpenStudy (anonymous):

because when you complete a square like this you need to both add and subtract

OpenStudy (anonymous):

@nphuongsun93 oooh ok now it makes sense lol.

OpenStudy (anonymous):

@LalanaG123 got the point i m trying to explain

OpenStudy (anonymous):

\[(a+b)^2 = a^2 + 2ab + b^2\]